OpenStudy (anonymous):

Marie and Tony went shopping together. Marie bought 2L of milk and 10 oranges for $3.44. Tony bought 3L of milk and 8 oranges for $3.69. What was the cost of one liter of milk? Of ine orange?

OpenStudy (ajprincess):

let us say that the cost of 1L of milk is x and cost of 1 orange is y. Then 2x+10y=3.44 3x+8y=3.69 2(x+5y)=2*1.72 x+5y=1.72 x=1.72-5y Substituting this in 3x+8y=3.69 3(1.72-5y)+8y=3.69 3*1.72-3*5y+8y=3.69 5.16-15y+8y=3.69 Find y from this and then plug in that value of y in x=1.72-5y to find x. Does that help?

OpenStudy (anonymous):

Yes wait i'll try it

OpenStudy (ajprincess):

ha k:)